OpenStudy (anonymous):

Determine a real conical form of the given matrix and give a change of coordinates matrix P that brings the matrix into this form. [ 0 -2 1 ] [ 2 2 -1 ] [ 0 -2 2 ]

OpenStudy (anonymous):

The matrix P I found was singular and therefore wrong (no P inverse exists). I can give you the diagonal of the matrix, canonical form and matrix P that I calculated but they could be wrong that is why I am waiting for someone to request them instead of just typing them.

OpenStudy (anonymous):

The matrix you listed isn't singular, right?

OpenStudy (anonymous):

It is the matrix given in the question and is not singular.

OpenStudy (anonymous):

Hmm, but the determinate....

OpenStudy (anonymous):

the determinate, calculated using a graphing calculator is 4.

OpenStudy (anonymous):

Doesn't the determinate have to be \(0\) for it to be singular?

OpenStudy (anonymous):

yes the matrix P that I found when I attempted the problem was [ 0.5 1 -0.5 ] [ 0 0.5 -0.5 ] [ 1 1 0 ] it is singular which does not fit with what P is suppose to be.

OpenStudy (anonymous):

Does that clear things up or should I try clarifying more?

OpenStudy (anonymous):

I'm very confused about what matrix is what right now.

OpenStudy (anonymous):

So the given matrix, let's call it \(A\) is some matrix. Then the matrix you typed out is \(B\). And \(PA=B\)? or something?

OpenStudy (anonymous):

Or maybe \[ PAP^{-1}=B \]

OpenStudy (anonymous):

yes but the matrix I most recently typed out is P. This is why I know it is wrong, the matrix P that I got has no inverse.

OpenStudy (anonymous):

I meant yes to the second formula you typed out

OpenStudy (anonymous):

I did find the value "B" in the formula you posted to be [ 2 0 0 ] [ 0 1 1 ] [ 0 -1 1 ]

OpenStudy (anonymous):

but that could be incorrect as well.

OpenStudy (anonymous):

B is suppose to be the real canonical form sorry I mispelt it in the original question

OpenStudy (anonymous):

\[ A = \begin{bmatrix} 0 & -2 & 1\\ 2 & 2 & -1\\ 1 & -2 & 2 \end{bmatrix} \]\[ B = \begin{bmatrix} 2 & 0 & 0\\ 0 & 1 & 1\\ 0 & -1 & 1 \end{bmatrix} \]

OpenStudy (anonymous):

So they give you \(A\) and \(B\)?

OpenStudy (anonymous):

No just A. I attempted to solve for the other matrices and was incorrect.

OpenStudy (anonymous):

What's a real canonical form? Can you use the diagonalization: \[ A=S^{-1}\Lambda S \]?

OpenStudy (anonymous):

Whoops, I mean \[ A = S\Lambda S^{-1} \]

OpenStudy (anonymous):

It uses imaginary numbers, and is similar but different to diagonalization. If you have not studied imaginary numbers I thank you for your time but you cannot help me.

OpenStudy (anonymous):

have you studied imaginary numbers?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

is that real canonical form? My textbook and professor have never referred to it as Jordan normal form.

OpenStudy (anonymous):

I haven't heard real canonical form mentioned anywhere, instead it goes to jordan canonical form.

OpenStudy (anonymous):

Different classes might have different names. This happens quite often.

OpenStudy (anonymous):

if a and b+ci are eignevalues of A then the real canonical form is [ a 0 0 ] [ 0 b c ] [ 0 -c b ]

OpenStudy (anonymous):

*eigenvalues

OpenStudy (anonymous):

does that fit the definition of Jordan canonical form?

OpenStudy (anonymous):

Very similar, but no, they seem different.

OpenStudy (anonymous):

Either way, you got eigen values of \(2, 1\pm i\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now I need help finding the matrix P that converts matrix A into its real conical form.

OpenStudy (anonymous):

You weren't taught any methodology for that?

OpenStudy (anonymous):

I am suppose to find eignvectors that correspond to the eigenvalues 2 and 1+i to convert them into the matrix P.

OpenStudy (anonymous):

I messed up my eigenvector corresponding to 1+i but I can't see where

OpenStudy (anonymous):

The only issue is that these vectors have imaginary components, so I suppose they are not the \(P\) you are looking for.

OpenStudy (anonymous):

yes they are I have another way to convert them into the matrix P that is similar to the way I stated you found the real canonical form.

OpenStudy (anonymous):

I thought the eigenvector corresponding to 1=i was [ 1 ] [-0.5] [0.5] + [-0.5]i [ 1 ] [ 0 ]

OpenStudy (anonymous):

sorry I meant corresponding to 1+i

OpenStudy (anonymous):

For \(1+i\) they have:\[ \begin{bmatrix} 0.5 \\ 0.5 \\ 1 \end{bmatrix} + \begin{bmatrix} 0.5 \\ -0.5 \\ 0 \end{bmatrix} i \]

OpenStudy (anonymous):

For \(1-i\) they have: \[ \begin{bmatrix} 0.5 \\ 0.5 \\ 1 \end{bmatrix} + \begin{bmatrix} -0.5 \\ 0.5 \\ 0 \end{bmatrix} i \]

OpenStudy (anonymous):

what do you get for \(P\) when you use these?

OpenStudy (anonymous):

I get the correct value of P which is (tried the operation in graphing calculator and it checked out) [ 0.5 0.5 0.5] [ 0 0.5 -0.5] [ 1 1 0 ]

OpenStudy (anonymous):

I have to use the basis for eigenvalue 2 to be [ 0.5 ] [ 0 ] [ 1 ]

OpenStudy (anonymous):

all they did was multiply it be two to get rid of the fraction but in order to get the correct value of P I cannot take that basis.

OpenStudy (anonymous):

How in the world are you entering a 3x3 matrix in the equation area?

OpenStudy (anonymous):

``` \begin{bmatrix} 11 & 12 & 13 \\ 21 & 22 & 23 \\ 31 & 32 & 33 \end{bmatrix} ```

OpenStudy (anonymous):

I type it all up without equation editor.

OpenStudy (anonymous):

Ooooooooooo fancy

OpenStudy (anonymous):

\begin{bmatrix}-1-i & -2 & 1 \\2 & 1-i & -1 \\0 & -2 & 1-i\end{bmatrix}

OpenStudy (anonymous):

I know the matrix P I originally had was not invertible that was how I knew it was incorrect.

OpenStudy (anonymous):

The matrix I just entered is what I started to try and row reduce

OpenStudy (anonymous):

then I went swapped the negative of row one with half of row two to get \begin{bmatrix}1 & 0.5(1-i) & -0.5 \\1+i & 2 & -1 \\0 & -2 & 1-i\end{bmatrix}

OpenStudy (anonymous):

Given the eigen vectors \(v_1, v_2, v_3\) how did you calculate \(P\)?

OpenStudy (anonymous):

then I subtracted (1+i)row one from row two to get \begin{bmatrix}1 & 0.5(1-i) & -0.5 \\0 & 1 & -0.5(1-i) \\0 & -2 & 1-i\end{bmatrix}

OpenStudy (anonymous):

Since I used the eigenvalue 1+i I use the eigenvectors corresponding to that eigenvalue and the whole number eigenvalue (2) If V1 is the eigenvector corresponding to eigenvalue 2 and V2 is the eigenvector corresponding to eigenvalue 1+i the I rewrite V2 and separate the real and imaginary components so that V2=W1+W2*i where W1 and W2 are vectors then P is the matrix with columns ( V1 W1 W2 )

OpenStudy (anonymous):

Ah, I see!

OpenStudy (anonymous):

thats why its so important that I messed up when calculating the eigenvector corresponding to 1+i

OpenStudy (anonymous):

Is there an easy way of finding \(P^{-1}\)?

OpenStudy (anonymous):

Do you just use Gauss-Jordan?

OpenStudy (anonymous):

not that I know of but luckily the question only asks for me to find P. I don't have to find its inverse

OpenStudy (anonymous):

I have no idea what that is but I used \[\det(A-\lambda I)=0\] to get the eigenvalues

OpenStudy (anonymous):

Man, the only trouble with linear algebra is that the methodology is so complicated it's easy to forget things.

OpenStudy (anonymous):

No Idea what what is?

OpenStudy (anonymous):

Oh, Gauss-Jordan method is just a way of inverting a matrix.

OpenStudy (anonymous):

then row reduced\[A-\lambda I\] to get the eigenvectors (usually you do this to get a basis for the eigenspace)

OpenStudy (anonymous):

Yeah, I have no problem with diagonalization with real eigen values.

OpenStudy (anonymous):

I'm just a bit rusty on complex and repeated eigenvalues.

OpenStudy (anonymous):

I don't either and getting the real canonical form is very similar but I seem to have messed up in my row reduction for eigenvalue 1+i and I don't know where thats why I started typing out my steps to see if you could spot where I went wrong in my row reduction.

OpenStudy (anonymous):

Solving for Null space is tricky enough... doing it with complex numbers must be really tricky.

OpenStudy (anonymous):

yes . . . and I think I may have just found my mistake one second

OpenStudy (anonymous):

I did now I have the correct eigenvector and therefore value of P. Thanks so much, I messed up in the last step of my row reduction which had two complex numbers multiplying. I did not calculate (1-i)(1-i) properly

OpenStudy (anonymous):

Thanks for sticking through it with me.

OpenStudy (anonymous):

no problem